OpenStudy (anonymous):

solve carlos scored 15 more points that john.if the product of the two points toal 154 how many points did each boy score

OpenStudy (anonymous):

x=15+y x*y=154 (15+y)*y=154 then just solve for y and plug that into the first eqn and solve for x

OpenStudy (radar):

Let x=john points x+15 = carlos points x(x+15) = 154 given x^2 + 15x =154 or x^2 + 15x -154=0 (x+22)(x-7)=0 x=7 x+15 = 22 carlos 22 points John 7 points